OpenStudy (anonymous):

kevin pichek 13 more apples that andy. If Kevin piched 30 apples, how many did Andi pick?

OpenStudy (texaschic101):

K = A + 13 K = 30 so sub 30 in for K in the first equation 30 = A + 13 30 - 13 = A 17 = A Andy picked 17 understand ?

OpenStudy (anonymous):

Got it!!!! thanks!!!=)